Without actually performing the long division, state whether the following rational numbers will have a terminating decimal expansion or a non-terminating repeating decimal expansion:
6/15

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

6/15= 2/5
2/5 = 0.4

Thanks